a woman is currently 21 years older than her daughter. Eleven years before the current time. the daughter was two-thirds as old as her mother. what is the current age of the mother?
Answer by mananth(16949)

a woman is currently 21 years older than her daughter. Eleven years before the current time. the daughter was two-thirds as old as her mother. what is the current age of the mother?
let daughter's age be x now
Mother's age now = (x+21)
11 years ago
daughter's age = (x-11)
Mother's age = (x+21)-11
(x-11)= (2/3)(x+10)
3(x-11)=2(x+10)
3x-33=2x+20
3x-2x=20+33
x=53
Daughter's age now = 53 years